OpenStudy (babynini):

Use the definition of the Definite Integral to evaluate the given integral. Show all steps and make sure to use the DEFINITION not the fundamental theorem of calculus! Make use of the identities.

OpenStudy (babynini):

I sort of have an idea of what i'm meant to do but not sure how to implement it o.o..ai.

OpenStudy (babynini):

@Miracrown are you available? :)

OpenStudy (babynini):

\[\int\limits_{1}^{3}x^3-x^2+2x-1=\lim_{h \rightarrow \infty}\sum_{i=1}^{n}f(x ^{*}_{i})\Delta x\]

OpenStudy (babynini):

Is the first step, I think. Although that is just a formula and now we have to find delta x

ganeshie8 (ganeshie8):

\[\int\limits_{1}^{3}x^3-x^2+2x-1=\lim_{\color{red}{n} \rightarrow \infty}\sum_{i=1}^{n}f(x ^{*}_{i})\Delta x\]

OpenStudy (babynini):

Delta x = 3-1/n = 2/n

ganeshie8 (ganeshie8):

yes

OpenStudy (babynini):

and using right end points we have a+i(Deltax) = 0+i(2/n) = 2i/n

ganeshie8 (ganeshie8):

good, but here a is not 0

OpenStudy (babynini):

whoops! o.o I had written it zero when I copied it.

OpenStudy (babynini):

(2i/n)+ 1 is that how I would write it?

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):

\(x_i^* = 1+(2i/n)\) \(f(x_i^*) = ?\)

OpenStudy (babynini):

eem i'm not sure D:

ganeshie8 (ganeshie8):

\(f(x)=x^3-x^2+2x-1\) \(f(x_i^*) = ?\)

OpenStudy (babynini):

oh i see

ganeshie8 (ganeshie8):

simply replace \(x\) by the value of \(x_i^*\)

OpenStudy (babynini):

Give me a moment :)

OpenStudy (babynini):

Can I even simplify the first one? the cubed one?

ganeshie8 (ganeshie8):

you may use (a+b)^3 formula

ganeshie8 (ganeshie8):

(a+b)^3 = a^3 + 3a^2b + 3ab^2 + b^3

OpenStudy (babynini):

hm okay. I have to go now but I shall check this in the morning and try to get my answer to you :) thanks!

ganeshie8 (ganeshie8):

good luck !

ganeshie8 (ganeshie8):

after simplifying, before evaluating the limit you should get http://www.wolframalpha.com/input/?i=%5Csum_%7Bi%3D1%7D%5En+%28%281%2Bi*2%2Fn%29%5E3-%281%2Bi*2%2Fn%29%5E2%2B2%281%2Bi*2%2Fn%29-1%29*2%2Fn